# Movable Antenna Empowered Near-Field Sensing via Antenna Position Optimization

### v1

@@ IV-A System Model
where $\widetilde{\text{CRB}}_{u}(\tilde{\boldsymbol{s}},\boldsymbol{\eta}_{\text{opt}})$, $\widetilde{\text{CRB}}_{v}(\tilde{\boldsymbol{s}},\boldsymbol{\eta}_{\text{opt}})$ and $\widetilde{\text{CRB}}_{r}(\tilde{\boldsymbol{s}},\boldsymbol{\eta}_{\text{opt}})$ denote the CRBs on the two AoAs and distance, respectively. It is observed that the objective function of (P2-3) contains even more complex terms w.r.t. the APM $\tilde{\boldsymbol{s}}$ compared with (P2-1) and (P2-2), as seen from (45), (46) and (47). As such, the discrete sampling-based algorithm is adopted again, with the main procedures the same as Algorithm 2.
@@ V Numerical Results
In this section, numerical results are presented to evaluate the performance of the proposed near-field MA sensing scheme for the estimation of target parameter(s) with both 1D and 2D MA arrays. Unless otherwise stated, the simulation parameters are as follows. The carrier frequency is set to 15 GHz so that the wavelength is $\lambda=0.02$ m. The minimum separation between adjacent MAs is set to $d=\lambda/2$. The average received signal-to-noise ratio (SNR) is defined as $P\lvert\beta\rvert^{2}/\sigma^{2}$. In the AoA estimation (Cases 1.1 and 2.1), we set the target distance as $r^{\star}=R_{RL}/4$. In the distance estimation, the target’s AoA is set as $\theta=45^{\circ}$, i.e., $u^{\star}=\cos\theta=0.71$, for the 1D MA array (Case 1.2); and we set $\theta=\phi=45^{\circ}$, i.e., $u^{\star}=\sin\theta\cos\phi=0.50$ and $v^{\star}=\cos\theta=0.71$, for the 2D MA array (Case 2.2). In the joint estimation of AoA and distance (Cases 1.2, 1.3, 2.2 and 2.3), we set $r_{\text{min}}=R_{FS}$ and $r_{\text{max}}=R_{RL}/2$, i.e., $r\in[R_{FS},R_{RL}/2]$. In Algorithms 1 and 2, we set the number of sampling points as M=10(N1)+1M=10(N-1)+1.
@@
To reveal more insights, we show the steering vector correlation under the three considered schemes in Figs. 9, 10 and 11, corresponding to Cases 1.1, 1.2, and 1.3, respectively. The steering vector correlation is defined as $R({\boldsymbol{\eta}}^{\prime})\triangleq\frac{1}{N^{2}}|\boldsymbol{\alpha}(\boldsymbol{x}^{\prime},\boldsymbol{\eta})^{\mathsf{H}}\boldsymbol{\alpha}(\boldsymbol{x}^{\prime},{\boldsymbol{\eta}}^{\prime})|^{2}$, where $\boldsymbol{\eta}$$^{\prime}$ denotes any feasible target parameter (vector), and $\boldsymbol{x}^{\prime}$ denotes the optimized MA position by any considered scheme. Evidently, it is desirable for the correlation function $R({\boldsymbol{\eta}}^{\prime})$ to approach a Dirac function, i.e., $\boldsymbol{R}({\boldsymbol{\eta}}^{\prime})\rightarrow\begin{cases}1,&{\boldsymbol{\eta}}^{\prime}=\boldsymbol{\eta},\\ 0,&{\boldsymbol{\eta}}^{\prime}\neq\boldsymbol{\eta},\end{cases}$ thereby achieving higher angular and spatial resolution while reducing ambiguity in parameter estimation. As observed from Figs. 9 and 10, the proposed scheme provides a narrower main lobe compared with Benchmarks 1 and 2. Moreover, it is observed from Fig. 9 that a grating lobe occurs at $0.23$ in Benchmark 2. This results in ambiguity in distinguishing the actual AoA $0.71$ from its false estimate at 0.23-0.23, leading to a large CRB for Benchmark 2. Additionally, by comparing Figs. 9 and 10, it is observed that the main lobe in the distance estimation is much broader than that in the AoA estimation, which indicates higher sensing accuracy in the angular domain compared with the distance domain. This is also revealed in Figs. 7(a) and 7(b), where the CRBs on the AoA estimation are significantly lower than those on the distance estimation. For the joint estimation, it is shown in Fig. 11(a) that the correlation function for the proposed scheme reaches its maximum at the location of the target with suppressed sidelobes in the angular domain, indicating enhanced resolution and reduced ambiguity. In contrast, Benchmark 1 yields a wider main lobe along the distance domain, thereby limiting its spatial resolution and leading to a larger CRB, as shown in Fig. 11(b). For Benchmark 2, Fig. 11(c) illustrates an additional undesired ridge-shaped pattern caused by the grating lobe at 0.23-0.23, which creates AoA estimation ambiguity and leads to a larger CRB compared with the proposed scheme.
@@
(a) Target distance of 5 m
(b) Target distance of 15 m
(c) Target distance of 30 m
@@
Note that we initialize the APM $\tilde{\boldsymbol{s}}$ as that in Benchmark 4 in Algorithm 2. The optimized MA positions for Cases 2.1, 2.2 and 2.3 by adopting Algorithm 2 are shown in Figs. 12(a)-(c), respectively. It is observed that there always exist MAs deployed on the corners of the square plane in the proposed scheme, thereby achieving the maximum aperture to improve the sensing resolution. Particularly, the array geometries in Cases 2.1 and 2.3 exhibit near symmetry along both dimensions, which achieves a balance between the estimation of $u$ and $v$, as evidenced in [ma2024movable]. Moreover, the optimized array geometry in Case 2.3 combines the key characteristics of both AoA and distance estimation. It preserves the MAs at the four corners in Case 2.1 and also the MAs around the central region in Case 2.2, thereby capturing the advantages of both configurations. Figs. 13(a)-(c) show the worst-case (sums) of the CRBs by different schemes versus the received SNR in Cases 2.1, 2.2 and 2.3, respectively. It is observed that the proposed scheme achieves significantly lower CRBs compared with Benchmarks 3 and 4. Similar trends to those observed in Fig. 7 for the 1D MA array are also observed in Fig. 13 for the 2D MA array. Thus, the detailed discussions are omitted.
Next, we examine the steering vector correlation for the optimized 2D MA array, i.e., $R({\boldsymbol{\eta}}^{\prime})$, in Case 2.3. With $v=0.71$, Figs. 14(a)-(c) show $R({\boldsymbol{\eta}}^{\prime})$ over the 2D angle-distance ($u$-$r$) domain for the target distances of 5 m, 15 m, and 30 m (all between $R_{RS}=0.54$ m and $R_{RS}=32.00$ m). Similar correlation patterns as those shown in Figs. 9, 10 and 11 are observed. Particularly, Fig. 14(a) reveals a sharply peaked main lobe at the true target location, highlighting the strong beam-focusing capability of the optimized array geometry for accurate target localization. However, as the target distance increases to 15 m and then to 30 m, undesired sidelobes along the AoA dimension gradually intensify, while the main lobe along the distance dimension becomes noticeably flatter, reducing its directivity and consequently degrading localization accuracy. This behavior is expected because larger target distances cause the spherical wavefronts to transition towards planar ones, thereby weakening near-field beamfocusing. Nevertheless, Fig. 14 also shows that the main lobe along the AoA direction remains sharply peaked even at large distances. This indicates that near-field target sensing remains largely comparable to its far-field counterpart in terms of AoA estimation.
@@
Finally, in Figs. 15 and 16, we compare the worst-case sums of the CRBs by different schemes in Case 2.3 versus the number of MAs $N$ and the side length of the movement region $A$, respectively. The received SNR is set to 10 dB. It is observed that the worst-case sums of the CRBs decrease with either $N$ or $A$, except for Benchmark 3 in Fig. 16, which remains unchanged with $A$, as its array geometry does not vary with the array aperture. Particularly, Fig. 15 shows that the proposed 2D MA array with only $N=9$ MAs achieves a 54.3% and a 98.4% reduction in the CRB compared with Benchmark 2 with $N=9$ MAs and Benchmark 3 with even $N=100$ MAs, respectively. In Fig. 16, even for a small-scale array with a side length of $A=0.2$ m, the proposed scheme is observed to lead to a 98.8% and an 18.5% decrease in the CRB compared with Benchmarks 3 and 4, respectively. These observations indicate that MAs offer a cost-effective solution for near-field target localization, as they can significantly reduce the number of required antennas compared with FPAs while even achieving superior localization accuracy.
@@ VI Conclusion
Based on [stoica1989music, Theorem 4.1], we can obtain the Fisher information of $u$ for estimating the AoA of the target as [^3] Note that unbiasedness and regularity of the estimators in all considered cases can be readily demonstrated. Therefore, the CRBs on both individual and joint estimation exist.
@@ Appendix C Proof of Theorem 2
$(N_{r}-1)(2N_{r}-1)d^{3}-2A^{3}dN_{r}^{2}(N_{r}-1)+\frac{N_{r}^{2}(N_{r}-1)(2N_{r}-1)A^{2}d^{2}}{3}-\frac{Ad^{3}}{3}N_{r}^{2}(N_{r}-1)^{2}(2N_{r}-1)\Big)^{2}\triangleq f(A,N).$ (70)
@@ Appendix D Proof of Corollary 2
$\overline{\text{CRB}}_{r}(\boldsymbol{x}^{\star})=\Big(\frac{2r^{2}_{\text{max}}}{1-{u^{\star}}^{2}}\Big)^{2}\cdot\frac{\kappa}{f(A,N)}.$ (71)
Based on (70), $f(A,N)$ is generally a quartic function w.r.t. $A$ as the term $\frac{N_{r}}{N}A^{4}$ dominates, and it is correlated with high-order terms w.r.t. $N$. Specifically, $f(A,N)$ increases with $A$ in the order of $\mathcal{O}(A^{4})$ for $A\geq(N-1)d$. It also increases with $N$ for $2\leq N\leq A/d+1$ but decreases with $d$ for $0\leq d\leq A/(N-1)$. Since $\overline{\text{CRB}}_{r}(\boldsymbol{x}^{\star})$ is inversely proportional to $f(A,N)$, we can infer that $\overline{\text{CRB}}_{r}(\boldsymbol{x}^{\star})$ decreases with $A$ for $A\geq(N-1)d$ in the order of $\mathcal{O}(A^{-4})$ and decreases with $N$ for $2\leq N\leq A/d+1$.
@@
By performing similar procedures to those in Appendix E, the CRBs on the two AoAs and distance estimation for the 2D MA array are given by (45), (46) and (47), respectively. This thus completes the derivations.

### v2

@@ IV-A System Model
$\quad+\widetilde{\text{CRB}}_{r}(\tilde{\boldsymbol{s}},\boldsymbol{\eta}_{\text{ed}})\Big)^{-1}$ (48a)
[…]
where $\widetilde{\text{CRB}}_{u}(\tilde{\boldsymbol{s}},\boldsymbol{\eta}_{\text{ed}})$, $\widetilde{\text{CRB}}_{v}(\tilde{\boldsymbol{s}},\boldsymbol{\eta}_{\text{ed}})$ and $\widetilde{\text{CRB}}_{r}(\tilde{\boldsymbol{s}},\boldsymbol{\eta}_{\text{ed}})$ denote the CRB on the two AoAs and distance, respectively. It is observed that the objective function of (P2-3) contains even more complex terms w.r.t. the APM $\tilde{\boldsymbol{s}}$ compared with (P2-1) and (P2-2), as seen from (45), (46) and (47). As such, the discrete sampling-based algorithm is adopted again, with the main procedures the same as Algorithm 2.
@@ V Numerical Results
In this section, numerical results are presented to evaluate the performance of the proposed near-field MA sensing scheme for the estimation of target parameter(s) with both 1D and 2D MA arrays. Unless otherwise stated, the simulation parameters are as follows. The carrier frequency is set to 15 GHz so that the wavelength is $\lambda=0.02$ m. The minimum separation between adjacent MAs is set to $d=\lambda/2$. The average received signal-to-noise ratio (SNR) is defined as $P\lvert\beta\rvert^{2}/\sigma^{2}$. For the 1D MA array, we set the ground-truth target AoAs in Cases 1.1 and 1.3 and the known target AoA in Case 1.2 as $\theta=45^{\circ}$, i.e., $u=\cos\theta=0.71$; and we set the ground-truth target distances in Cases 1.2 and 1.3 and the known target distance in Case 1.1 as $r=R_{RL}/4$. For the 2D MA array, we set the ground-truth target AoAs in Cases 2.1 and 2.3 and the known target AoA in Case 2.2 as $\theta=\phi=45^{\circ}$, i.e., $u=\sin\theta\cos\phi=0.50$ and $v=\cos\theta=0.71$; and we set the ground-truth target distances in Cases 2.2 and 2.3 and the known target distance in Case 2.1 as $r=R_{RL}/4$. Moreover, in Cases 1.2, 1.3, 2.2 and 2.3, we set $r_{\text{min}}=R_{FS}$ and $r_{\text{max}}=R_{RL}/2$, i.e., $r\in[R_{FS},R_{RL}/2]$. In Cases 1.1 and 1.3, we set $\theta_{\text{min}}=\pi/10$ with $u_{\text{max}}=0.95$; while in Cases 2.1 and 2.3, we set $\theta_{\text{min}}=\phi_{\text{min}}=\pi/10$ with $u_{\text{max}}=v_{\text{max}}=0.95$. In Algorithms 1 and 2, we set the number of sampling points as M=10(N1)+1M=10(N-1)+1.
@@
To reveal more insights, we show the steering vector correlation under the three considered schemes in Figs. 8, 9 and 10, corresponding to Cases 1.1, 1.2, and 1.3, respectively. The steering vector correlation is defined as $R({\boldsymbol{\eta}}^{\prime})\triangleq\frac{1}{N^{2}}|\boldsymbol{\alpha}(\boldsymbol{x}^{\prime},\boldsymbol{\eta})^{\mathsf{H}}\boldsymbol{\alpha}(\boldsymbol{x}^{\prime},{\boldsymbol{\eta}}^{\prime})|^{2}$, where $\boldsymbol{\eta}$$^{\prime}$ denotes any feasible target parameter (vector), and $\boldsymbol{x}^{\prime}$ denotes the optimized APV by any considered scheme. Evidently, it is desirable for the correlation function $R({\boldsymbol{\eta}}^{\prime})$ to approach a Dirac function, i.e., $\boldsymbol{R}({\boldsymbol{\eta}}^{\prime})\rightarrow\begin{cases}1,&{\boldsymbol{\eta}}^{\prime}=\boldsymbol{\eta},\\ 0,&{\boldsymbol{\eta}}^{\prime}\neq\boldsymbol{\eta},\end{cases}$ thereby achieving higher angular and spatial resolution while reducing ambiguity in parameter estimation. As observed from Figs. 8 and 9, the proposed scheme provides a narrower main lobe compared with Benchmarks 1 and 2. Moreover, it is observed from Fig. 8 that a grating lobe occurs at 0.23-0.23 in Benchmark 2. This results in ambiguity in distinguishing the actual AoA $0.71$ from its false estimate at 0.23-0.23, leading to a large CRB for Benchmark 2. Additionally, by comparing Figs. 8 and 9, it is observed that the main lobe in the distance estimation is much broader than that in the AoA estimation, which indicates higher sensing accuracy in the angular domain compared with the distance domain. This is also revealed in Figs. 6(a) and 6(b), where the CRBs on the AoA estimation are significantly lower than those on the distance estimation. For the joint estimation, we map $R({\boldsymbol{\eta}}^{\prime})$ onto the $x$-$y$ domain with $(x,y)=(ru,r\sqrt{1-u^{2}})$ and present the results for the proposed scheme, Benchmark 1, and Benchmark 2 in Figs. 10(a)-(c), respectively. It is observed from Fig. 10(a) that the correlation function for the proposed scheme reaches its maximum at the location of the target with suppressed sidelobes in the angular domain, indicating enhanced resolution and reduced ambiguity. In contrast, Benchmark 1 yields a wider main lobe along the distance domain, thereby limiting its spatial resolution and leading to a larger CRB, as shown in Fig. 10(b). For Benchmark 2, Fig. 10(c) illustrates an additional undesired ridge-shaped pattern caused by the grating lobe at 0.23-0.23, which creates AoA estimation ambiguity and leads to a larger CRB compared with the proposed scheme.
@@
(a) Target distance of 3 m
(b) Target distance of 8 m
(c) Target distance of 15 m
@@
Note that we initialize the APM $\tilde{\boldsymbol{s}}$ as that in Benchmark 4 in Algorithm 2. The optimized MA positions for Cases 2.1, 2.2 and 2.3 by adopting Algorithm 2 are shown in Figs. 11(a)-(c), respectively. It is observed that there always exist MAs deployed on the corners of the square plane in the proposed scheme, thereby achieving the maximum aperture to improve the sensing resolution. Particularly, Figs. 11(a) and (c) show that the array geometries in Cases 2.1 and 2.3 exhibit near symmetry along both dimensions, which achieves a balance between the estimation of $u$ and $v$, as evidenced in [ma2024movable]. Moreover, the optimized array geometry in Case 2.3 combines the key characteristics of both AoA and distance estimation, as shown in Fig. 11(c). It preserves the MAs at the four corners in Case 2.1 and also the MAs around the central region in Case 2.2, thereby capturing the advantages of both configurations. Figs. 12(a)-(c) show the worst-case (sums) of the CRBs by different schemes versus the received SNR in Cases 2.1, 2.2 and 2.3, respectively. It is observed that the proposed scheme achieves significantly lower CRBs compared with Benchmarks 3 and 4. Similar trends to those observed in Fig. 6 for the 1D MA array are also observed in Fig. 12 for the 2D MA array. Thus, the detailed discussions are omitted.
Next, we examine the steering vector correlation for the optimized 2D MA array, i.e., $R({\boldsymbol{\eta}}^{\prime})$, in Case 2.3, by mapping it onto the $x$-$y$ domain with $(x,y)=(ru,rv)$. With $v=0.71$, Figs. 13(a)-(c) show $R({\boldsymbol{\eta}}^{\prime})$ over the $x$-$y$ domain for the target distances of 3 m, 8 m and 15 m (all between $r_{\text{min}}=0.54$ m and $r_{\text{max}}=16.00$ m). Similar correlation patterns as those shown in Figs. 8, 9 and 10 are observed. Particularly, Fig. 13(a) reveals a sharply peaked main lobe at the true target location, highlighting the strong beamfocusing capability of the optimized array geometry for accurate target localization. However, as the target distance increases to 8 m and then to 15 m, undesired sidelobes along the AoA dimension gradually intensify, while the main lobe along the distance dimension becomes noticeably flatter, reducing its directivity and consequently degrading localization accuracy. This behavior is expected because larger target distances cause the spherical wavefronts to gradually transition towards planar ones, thereby weakening near-field beamfocusing. Nevertheless, Fig. 13 also shows that the main lobe along the AoA direction remains sharply peaked even at large distances. This indicates that near-field target sensing remains largely comparable to its far-field counterpart in terms of AoA estimation.
@@
Finally, in Figs. 14 and 15, we compare the worst-case sums of the CRBs by different schemes in Case 2.3 versus the number of MAs $N$ and the side length of the movement region $A$, respectively. The received SNR is set to 10 dB. It is observed that the worst-case sums of the CRBs decrease with either $N$ or $A$, except for Benchmark 3 in Fig. 15, which remains unchanged with $A$, as its array geometry does not vary with the array aperture. Particularly, Fig. 14 shows that the proposed 2D MA array with only $N=9$ MAs achieves a 24.5% and a 98.8% reduction in the CRB compared with Benchmark 4 with $N=9$ MAs and Benchmark 3 with even $N=100$ MAs, respectively. In Fig. 15, even for a small-scale array with a side length of $A=0.2$ m, the proposed scheme is observed to lead to a 99.2% and a 45.5% decrease in the CRB compared with Benchmarks 3 and 4, respectively. These observations indicate that MAs offer a cost-effective solution for near-field target localization, as they can significantly reduce the number of required antennas compared with FPAs while even achieving superior localization accuracy.
@@ VI Conclusion
Based on [stoica1989music, Theorem 4.1], we can obtain the Fisher information of $u$ for estimating the AoA of the target as [^3] Note that unbiasedness and regularity of the estimators in all considered cases can be readily demonstrated. Therefore, the CRBs on the estimation MSEs for both individual and joint estimation exist.
@@ Appendix C Proof of Theorem 2
$(N_{r}-1)(2N_{r}-1)d^{3}-2A^{3}dN_{r}^{2}(N_{r}-1)+\frac{N_{r}^{2}(N_{r}-1)(2N_{r}-1)A^{2}d^{2}}{3}-\frac{Ad^{3}}{3}N_{r}^{2}(N_{r}-1)^{2}(2N_{r}-1)\Big)^{2}\triangleq f(A,N,d).$ (70)
@@ Appendix D Proof of Corollary 2
$\text{CRB}_{r}(\boldsymbol{x}^{\star},r_{\text{max}})=\Big(\frac{2r^{2}_{\text{max}}}{1-{u^{\star}}^{2}}\Big)^{2}\cdot\frac{\kappa}{f(A,N,d)}.$ (71)
Based on (70), $f(A,N,d)$ is generally a quartic function w.r.t. $A$ as the term $\frac{N_{r}}{N}A^{4}$ dominates, and it is correlated with high-order terms w.r.t. $N$. Specifically, $f(A,N,d)$ increases with $A$ in the order of $\mathcal{O}(A^{4})$ for $A\geq(N-1)d$. It also increases with $N$ for $2\leq N\leq A/d+1$ but decreases with $d$ for $0\leq d\leq A/(N-1)$. Since $\text{CRB}_{r}(\boldsymbol{x}^{\star},r_{\text{max}})$ is inversely proportional to $f(A,N,d)$, we can infer that $\text{CRB}_{r}(\boldsymbol{x}^{\star},r_{\text{max}})$ decreases with $A$ for $A\geq(N-1)d$ in the order of $\mathcal{O}(A^{-4})$ and decreases with $N$ for $2\leq N\leq A/d+1$.
@@
By performing similar procedures to those in Appendix E, the CRBs for the two AoAs and the distance are obtained as (45), (46), and (47), respectively. This thus completes the derivations.